### v2

# DocMath-Eval: Evaluating Math Reasoning Capabilities of LLMs in Understanding Long and Specialized Documents

Abstract
Recent LLMs have demonstrated remarkable performance in solving exam-like math word problems. However, the degree to which these numerical reasoning skills are effective in real-world scenarios, particularly in expert domains, is still largely unexplored. This paper introduces DocMath-Eval, a comprehensive benchmark specifically designed to evaluate the numerical reasoning capabilities of LLMs in the context of understanding and analyzing specialized documents containing both text and tables. We evaluate a wide spectrum of 48 LLMs with Chain-of-Thought and Program-of-Thought prompting methods, aiming to comprehensively assess the capabilities and limitations of existing LLMs in DocMath-Eval. We found that even the current best-performing system (i.e., GPT-4o) still significantly lags behind human experts in solving complex numerical reasoning problems grounded in long contexts. We believe that DocMath-Eval can serve as a valuable benchmark for evaluating LLMs’ capabilities in solving challenging numerical reasoning problems within expert domains.
DocMath-Eval: Evaluating Math Reasoning Capabilities of LLMs
in Understanding Long and Specialized Documents
Yilun Zhao††thanks: Equal Contributions. 1 Yitao Long Hongjun Liu2 Ryo Kamoi3 Linyong Nan1
Lyuhao Chen4 Yixin Liu1 Xiangru Tang1 Rui Zhang3 Arman Cohan1,5
1Yale University 2New York University 3Penn State University
4Carnegie Mellon University 5Allen Institute for AI
1 Introduction
Recent advancements in large language models (LLMs) have attracted significant attention due to their capabilities in solving a broad range of tasks OpenAI (2022, 2023); Touvron et al. (2023), including math word problems (MWPs) commonly found in academic exams Wang et al. (2017); Miao et al. (2020); Amini et al. (2019); Cobbe et al. (2021); Hendrycks et al. (2021); Cobbe et al. (2021); Lu et al. (2023); Chen et al. (2023b). These MWPs vary from basic arithmetic to advanced algebra, showcasing LLMs’ proficiency in numerical reasoning — a crucial skill for interpreting and manipulating numerical data across various contexts. Despite this progress, there is still a significant gap in understanding the practicality of LLMs’ numerical reasoning in real-world scenarios, particularly in specialized fields such as finance, medicine, and science. As illustrated in Figure 1, these expert domains necessitate LLMs to interpret complex, domain-specific documents, applying numerical reasoning to complex problem-solving Chen et al. (2021); Zhu et al. (2021); Zhao et al. (2022); Li et al. (2022b). Recognizing this gap, our research focuses on the finance domain Li et al. (2022a); Wu et al. (2023a); Yang et al. (2023); Callanan et al. (2023); Xie et al. (2024). The finance industry often deals with lengthy and data-intensive documents that demand advanced numerical reasoning skills for accurate analysis and decision-making.
We introduce DocMath-Eval, a comprehensive and standardized benchmark that systematically evaluates the numerical reasoning capabilities of LLMs in understanding and interpreting specialized documents containing both textual and tabular data. DocMath-Eval encompasses four evaluation sets, each with varying levels of difficulty in numerical reasoning and document understanding. Specifically, We construct a new evaluation set, DM, from scratch, to examine the LLM’s capabilities in performing complex numerical reasoning over extreme long documents containing multiple tables. We also adapt and re-annotate four existing finance QA benchmarks to develop three additional, less challenging evaluation sets: 1) DM based on TAT-QA Zhu et al. (2021) and FinQA Chen et al. (2021), necessitates simple numerical reasoning over short document with one table; 2) DM based on MultiHiertt Zhao et al. (2022), necessitates simple numerical reasoning over long document with multiple tables; and 3) DM based on TAT-HQA Li et al. (2022b), necessitates complex numerical reasoning over short document with one table.
We evaluate a wide spectrum of open- and closed-source LLMs, specifically, 48 models from 17 organizations. Two prompting methods, Chain-of-Thought (CoT) Wei et al. (2022) and Program-of-Thought (PoT) Chen et al. (2023a), are adopted for experiments. Our experimental results indicate that while the existing best-performing LLM on average (i.e., GPT-4o) can achieve high performance in simple settings (e.g., DM), it still falls short of human experts in more challenging ones, i.e.,, DM. Moreover, Claude-3.5-Sonnet outperforms other LLMs, achieving an accuracy of 40.0% on the DM set when applying CoT prompting. However, it still lags far behind human expert performance, which stands at 76%. This significant gap between LLMs and human experts underscores the challenges presented by DocMath-Eval. It highlights the need for further advancements in adapting LLMs’ numerical reasoning capabilities for practical application in real-world expert domains.
We conclude our main contributions as follows:
We introduce DocMath-Eval, a comprehensive benchmark designed to systematically evaluate LLMs’ numerical reasoning ability to understand and interpret long and specialized documents. This includes a newly developed, challenging evaluation set and three adapted evaluation sets for varying difficulty levels.
We conduct an extensive evaluation encompassing a wide range of LLMs, including those specialized in math and coding. We also incorporate different prompting methods (i.e., CoT and PoT) to comprehensively assess the capabilities and limitations of existing LLMs in our task.
Our experimental results reveal a noticeable performance gap compared to human experts in more complex scenarios (i.e., problems requiring complex numerical reasoning over long documents). This highlights the limitations of current LLMs in complex real-world applications and the need for continued advancements.
Property (Median/Avg) | DM | DM | DM | DM (new) |
Data Source | TAT-QA Zhu et al. (2021) | MultiHiertt | TAT-HQA | expert annotated |
FinQA Chen et al. (2021) | Zhao et al. (2022) | Li et al. (2022b) | from scratch | |
Question Length | 19 / 20.0 | 21 / 21.6 | 29 / 30.1 | 34 / 37.7 |
# Sentences in Text | 14 / 16.9 | 64 / 66.9 | 6 / 7.8 | 535 / 752.3 |
# Words in Text | 504 / 506.6 | 2,216 / 2,334.0 | 251 / 314.2 | 25,149 / 34,589.0 |
# Table | 1 / 1.0 | 4 / 3.9 | 1 / 1.0 | 46 / 72.5 |
# Rows per Table | 6 / 7.0 | 9 / 11.6 | 7 / 8.2 | 3 / 7.5 |
# Columns per Table | 5 / 4.7 | 4 / 4.5 | 5 / 5.0 | 3 / 3.1 |
# Text Evidence | 0 / 0.4 | 1 / 0.9 | 0 / 0.4 | 1 / 1.0 |
# Table Evidence | 1 / 0.9 | 1 / 1.1 | 1 / 1.0 | 1 / 1.0 |
% Questions Table Evidence | 92.9% | 86.4% | 97.8% | 76.3% |
# Math Operations in Python Solution | 2 / 2.1 | 2 / 2.3 | 2 / 2.3 | 4 / 4.9 |
# Code Lines in Python Solution | 5 / 5.3 | 6 / 5.9 | 5 / 5.3 | 8 / 8.2 |
# Comment Lines in Python Solution | 2 / 2.0 | 2 / 2.0 | 2 / 2.0 | 2 / 3.4 |
Development set | 200 | 100 | 200 | 300 |
Test set | 800 | 400 | 800 | 1,200 |
Total Size | 1,000 | 500 | 1,000 | 1,500 |
2 Related Work
Math Word Problems
The research community has shown significant interest in the vital role of numerical reasoning skills in LLMs. These skills are vital for models to effectively engage in complex problem-solving. To this end, a wide variety of MWP datasets have been proposed in recent years Hosseini et al. (2014); Koncel-Kedziorski et al. (2016); Wang et al. (2017); Ling et al. (2017); Cobbe et al. (2021). More challenging datasets have recently been introduced to enhance diversity Miao et al. (2020), difficulty Chen et al. (2023b); Hendrycks et al. (2021), and adversarial robustness Patel et al. (2021). However, existing MWP datasets predominantly focus on problems akin to academic exams, with a limited emphasis on real-world scenarios. Addressing this gap, our paper introduces a novel and comprehensive benchmark designed to evaluate LLMs’ abilities in understanding and interpreting long and specialized documents through numerical reasoning.
Numerical Reasoning over Documents
Numerical reasoning over documents requires models to have a deep understanding of context and the ability to derive answers through numerical reasoning Dua et al. (2019). Applying these models in the finance domain Xie et al. (2023); Wu et al. (2023a); Yang et al. (2023) presents additional challenges in terms of interpreting hybrid data Zhu et al. (2021) and utilizing domain-specific expertise Chen et al. (2021); Zhao et al. (2023a). Numerous datasets focusing on numerical reasoning over specialized documents have been proposed recently. Two notable benchmarks are TAT-QA Zhu et al. (2021) and FinQA Chen et al. (2021), which represent pioneering efforts in studying numerical reasoning in finance, particularly requiring the fusion of tabular and textual content. Building upon TAT-QA, a more challenging dataset named TAT-HQA Li et al. (2022b) was developed, focusing on counterfactual questions in relation to the provided context. Additionally, MultiHiertt Zhao et al. (2022) focuses on numerical reasoning over longer financial documents containing multiple tables. However, as illustrated in Table 1, these four datasets focus on less challenging scenarios, where either simple numerical reasoning (e.g., calculating the increasing rate or average value) is sufficient, or the input context is short. Furthermore, there is a lack of a standardized benchmark for systematically evaluating models’ performance across varying difficulty levels in terms of numerical reasoning and document understanding.
3 DocMath-Eval
In this section, we first offer a formal definition of the DocMath-Eval task. We then explain the rationale and methodology for adopting Python program as the standardized solution format for DocMath-Eval. Subsequently, we detail the data annotation process used to construct the challenging DM evaluation set, as well as the data re-annotation process for compiling the other three evaluation sets. Table 7 in the Appendix presents the profiles of the seven annotators involved. Finally, we present human-level performance on each evaluation set in DocMath-Eval.
3.1 Task Formulation
We formally define the task of DocMath-Eval in the context of LLMs as follows: Presented with a numerical reasoning question and a financial document consisting of textual contents and structured tables , the task is to generate the numeric-value answer :
To obtain the best candidate answer , we use greedy decoding in all our LLM evaluations.
3.2 Solution Format Standardization
We observe that existing finance QA datasets feature solutions in various formats. Specifically, TAT-QA Zhu et al. (2021) and TAT-HQA Li et al. (2022b) utilize text, while MultiHiertt Zhao et al. (2022) employs mathematical expressions, such as 100/3, and FinQA Chen et al. (2021) uses math programs, such as divide(100,3), for solution annotations. This diversity in annotation formats hinders the development of a unified evaluation framework to assess LLM performance across different benchmarks. Additionally, solutions in text format often lack the precision and unambiguity necessary for computational problem-solving; and solutions in mathematical equations or math programs are less descriptive, with the semantic meaning of the equations sometimes being unclear.
To overcome the aforementioned limitations, in DocMath-Eval, we represent solutions using Python programs. Such a unified Python program format supports a standardized and effective evaluation framework for LLM assessment. Specifically, annotators are required to first define variables at the beginning of the Python function, starting with “def solution():”. These variables correspond to the key elements or quantities mentioned in the question or question-relevant content in the documents. Annotators are instructed to assign meaningful names that clearly represent each element. They then write a sequence of Python statements that logically solve the problem, step by step. Additionally, annotators receive a bonus for writing detailed comments, thereby enhancing the code’s readability and understandability. To ensure the accuracy and functionality of the solutions, our annotation interface automatically executes the Python function. This execution checks that the return type of the answer is either a float or an int and verifies that there are no execution errors.
3.3 Data Re-Annotation From Public Datasets
We re-annotate four existing datasets and incorporate them into DocMath-Eval. Specifically, we re-annotate TAT-QA Zhu et al. (2021) and FinQA Chen et al. (2021) for DM, MultiHiertt Zhao et al. (2022) for DM, and TAT-HQA Li et al. (2022b) for DM.
Question Validation and Re-annotation
We instruct the annotators to identify and remove questions with incorrect annotations or those whose answers are not numerical. Annotators are then asked to enhance each question by adding a scale descriptor to ensure clarity and specificity. For example, "Question: What is the average payment volume per transaction for American Express? (in billions)". They were also asked to correct any identified errors in the original questions.
Solution Validation and Re-annotation
As outlined in Section 3.2, we require annotators to rewrite the original solutions into a unified Python format, standardizing variable names and adding comments to enhance the readability of the solutions. Regarding the supporting evidence annotation, we initially convert the original evidence annotations to our format. We then highlight these evidences in the annotation interface, and direct annotators to verify their correctness.
3.4 Data Annotation From Scratch
In real-world scenarios, financial professionals typically need to handle documents spanning tens of pages, along with problems that require more complex numerical reasoning combined with financial knowledge. However, as previously discussed, existing benchmarks Zhu et al. (2021); Chen et al. (2021); Zhao et al. (2022); Li et al. (2022b) focus on less challenging scenarios, where either simple numerical reasoning is sufficient, or the input context is short. To bridge this gap, we have developed a new, challenging evaluation set, DM, from scratch. This set focuses on settings that more closely align with real-world scenarios, where models are required to perform complex numerical reasoning over long financial documents for problem solving. The annotation process is as follows:
Source Document Collection
Following previous work Zhu et al. (2021); Chen et al. (2021); Zhao et al. (2022), we use the quarterly (i.e., Form 10-Q) and annual reports (i.e., Form 10-K) of companies as our source documents, which are publicly available at the open-source database111https://www.sec.gov/edgar/search/ of U.S. Securities and Exchange Commission. After collecting all the source documents, we utilize a commercial API222https://sec-api.io/ to extract their textual and tabular content. Subsequently, we apply a heuristic-based method to preprocess these two formats of content. The preprocessed documents are then passed to expert annotators for question annotation.
Data Annotation
Given a financial document, annotators are first required to briefly read its content and determine the data points to be used in the question. They must then compose the question and highlight the selected paragraphs or tables as evidence supporting it. Finally, the annotators are required to write down the solution to the question in Python program format, as discussed in Section 3.2. We set up a bonus payment system for complex annotations that involve difficult document comprehension and numerical reasoning. Specifically, to increase the difficulty of document understanding, we award bonuses to annotators for questions that necessitate information from: 1) multiple tables, 2) multiple sections, or 3) a combination of tables and textual content. To enhance the challenge in numerical reasoning, we provide bonuses for questions requiring financial expertise or involving complex mathematical operations. If such annotations are validated during the quality validation stage, a bonus payment will be added.
Quality Validation
We implement a comprehensive quality validation protocol to ensure that each annotated example meets the required standards. For every question annotation, we assign it to another annotator, recognized for their high performance in annotation, to verify its accuracy. This process involves manually locating the question-relevant evidence in the documents using our retrieval-based search toolkits. They then compare this evidence with the original annotations and correct any errors found. Additionally, validators are tasked with confirming the accuracy of the annotated solutions. We offer bonus payments to annotators for identifying erroneous annotations. Ultimately, 232 of the annotated questions are flagged as erroneous and are subsequently revised. Appendix A in the Appendix presents the human evaluation scores and inter-evaluator agreements for a subset of 200 sampled examples. DocMath-Eval exhibits superior annotation quality and a high degree of inter-annotator agreement.
3.5 Expert-level Performance Evaluation
To provide a rough but informative estimate of the performance of domain-experts on each of DocMath-Eval sets, we invite two professionals with Chartered Financial Analyst licenses for evaluation. Regarding human expert performance on DM and DM, we report the same results as those in the original papers, with accuracy of 91% and 87%, respectively. For DM and DM, We randomly sample 25 examples from each set, asking the expert evaluators to answer the questions individually within a four-hour period. They achieve accuracy of 88% and 80% on DM (average 84%); and accuracy of 72% and 80% on DM (average 76%).
3.6 Dataset Release
Table 1 describes the basic statistics of four developed evaluation sets. DocMath-Eval contains a total of 4,000 questions with high-quality annotations, featuring varying difficulty levels in numerical reasoning and document understanding. We randomly divided the dataset into two splits: testmini and test. The testmini set contains 800 examples and is intended for model development validation. The test set comprises the remaining 3,200 examples and is designed for standard evaluation. To prevent data contamination Deng et al. (2024), the features relevant to the ground truth for the test set are not publicly released. Instead, we developed and maintain an online evaluation platform, allowing researchers to evaluate models and participate in a leaderboard.
4 Experiment Setup
This section discusses the experiment setup, including the evaluated LLMs, prompting methods, and our implementation details.
4.1 Large Language Models
Our goal is to investigate the capabilities of current state-of-the-art LLMs on DocMath-Eval to better understand their strengths and limitations. To this end, we evaluate a wide range of models:
General: GPT-3.5&4 OpenAI (2022, 2023, 2024), Gemini-1.5 Gemini (2024), Claude-3&3.5 Anthropic (2024), Llama-2&3&3.1 Touvron et al. (2023); AI@Meta (2024), Mistral Jiang et al. (2023), Phi-3 Abdin et al. (2024), Gemma-1&2 Team et al. (2024), WizardLM-2 Xu et al. (2023), Yi-1.5 01.AI (2023), Qwen-2 Bai et al. (2023), Command R+ Cohere (2024b), Aya Cohere (2024a), and GLM-4 GLM et al. (2024).
We select the most recent checkpoint available as of August 1, 2024. The details of each evaluated model, including the exact model version we used, are presented in Appendix A in the Appendix.
4.2 Prompting Methods
Following recent LLM reasoning benchmark works Lu et al. (2024); Chen et al. (2023b), we evaluate two established prompting methods:
Chain-of-Thought
Program-of-Thought
Different from CoT, the PoT method Chen et al. (2023a) disentangles computation from the reasoning process by prompting the LLMs to generate a structured program to represent the reasoning process. The final answer is then derived by executing the generated program with an external calculator. Figure 3 in Appendix presents the PoT prompt used in our experiment.
Model | Size | Notes | DM | DM | DM | DM | Avg. Acc | |||||||||
PoT | CoT | PoT | CoT | PoT | CoT | PoT | CoT | PoT | CoT | |||||||
Human Expert | 91.0 | 87.0 | 84.0 | 76.0 | ||||||||||||
Proprietary LLMs | ||||||||||||||||
GPT-4o | 84.0 | 86.0 | 69.5 | 76.5 | 56.0 | 64.0 | 41.0 | 36.7 | 60.8 | 62.4 | ||||||
GPT-4-Turbo | 85.5 | 82.5 | 80.0 | 81.0 | 56.0 | 53.0 | 38.7 | 38.3 | 62.9 | 61.9 | ||||||
Claude-3-Opus | 80.5 | 79.5 | 73.5 | 77.5 | 51.0 | 61.0 | 42.0 | 39.7 | 60.6 | 61.8 | ||||||
Claude-3.5-Sonnet | 78.0 | 77.0 | 76.0 | 69.5 | 54.0 | 61.0 | 44.0 | 40.0 | 61.8 | 59.2 | ||||||
Claude-3-Sonnet | 82.5 | 80.0 | 80.5 | 73.0 | 55.0 | 56.0 | 40.3 | 35.3 | 62.7 | 58.5 | ||||||
Gemini-1.5-Flash | 85.0 | 78.0 | 78.5 | 69.5 | 55.0 | 46.0 | 40.0 | 31.7 | 62.8 | 54.5 | ||||||
Gemini-1.5-Pro | 85.5 | 80.5 | 80.0 | 58.0 | 58.0 | 55.0 | 40.3 | 30.0 | 63.7 | 52.8 | ||||||
Claude-3-Haiku | 74.5 | 79.0 | 71.5 | 58.5 | 55.0 | 50.0 | 36.7 | 31.7 | 57.1 | 52.5 | ||||||
GPT-4o-Mini | 88.5 | 69.5 | 77.0 | 69.5 | 53.0 | 56.0 | 38.7 | 28.0 | 62.5 | 52.2 | ||||||
GPT-3.5-Turbo | 71.0 | 60.5 | 52.5 | 39.0 | 41.0 | 28.0 | 28.7 | 15.0 | 46.8 | 34.0 | ||||||
Open-source LLMs | ||||||||||||||||
DeepSeek-V2 | 236B | MoE | 87.0 | 82.0 | 75.5 | 69.5 | 61.0 | 56.0 | 43.0 | 39.7 | 64.4 | 59.8 | ||||
Mistral-Large | 123B | 85.0 | 83.5 | 76.5 | 81.0 | 56.0 | 55.0 | 41.0 | 31.3 | 62.8 | 59.7 | |||||
DeepSeek-Coder-V2 | 236B | Code | 85.0 | 79.0 | 78.0 | 66.5 | 56.0 | 54.0 | 41.0 | 37.7 | 63.1 | 57.3 | ||||
Llama-3.1 | 70B | 74.5 | 76.5 | 68.0 | 71.0 | 53.0 | 50.0 | 34.7 | 29.3 | 55.3 | 54.1 | |||||
Qwen2 | 72B | 26.5 | 74.0 | 24.5 | 72.5 | 8.0 | 45.0 | 7.0 | 27.0 | 16.4 | 52.4 | |||||
Llama-3 | 70B | 84.5 | 73.5 | 64.0 | 63.5 | 52.0 | 42.0 | 41.0 | 28.3 | 59.0 | 50.1 | |||||
Mixtral-8x22B | 141B | MoE | 30.0 | 74.0 | 21.5 | 57.0 | 25.0 | 47.0 | 14.7 | 24.0 | 21.5 | 47.6 | ||||
Gemma-2 | 9B | 79.0 | 66.5 | 65.0 | 54.5 | 50.0 | 39.0 | 24.3 | 17.7 | 51.4 | 41.8 | |||||
DeepSeek-Coder-V2-Lite | 16B | Code | 66.0 | 67.5 | 51.0 | 53.5 | 27.0 | 30.0 | 22.0 | 20.3 | 40.9 | 41.6 | ||||
WizardLM-2 | 141B | MoE | 62.5 | 60.5 | 56.5 | 55.5 | 25.0 | 34.0 | 17.7 | 18.0 | 39.5 | 40.0 | ||||
C4AI Command R+ | 104B | 35.5 | 65.5 | 39.0 | 51.0 | 19.0 | 31.0 | 8.7 | 18.3 | 24.3 | 39.9 | |||||
Yi-1.5 | 9B | 18.0 | 68.5 | 24.5 | 56.0 | 2.0 | 14.0 | 4.0 | 14.0 | 12.4 | 38.1 | |||||
Yi-1.5 | 34B | 0.5 | 64.5 | 1.0 | 53.0 | 0.0 | 14.0 | 0.0 | 15.3 | 0.4 | 36.9 | |||||
Mistral-Nemo | 12B | 52.5 | 59.5 | 37.5 | 44.0 | 28.0 | 37.0 | 15.3 | 16.7 | 31.7 | 36.8 | |||||
Llama-3.1 | 8B | 62.0 | 60.0 | 44.0 | 42.5 | 32.0 | 33.0 | 19.0 | 14.3 | 37.6 | 35.1 | |||||
DBRX | 132B | MoE | 41.0 | 57.0 | 29.5 | 43.0 | 32.0 | 30.0 | 12.0 | 16.3 | 26.1 | 34.9 | ||||
Codestral | 22B | Code | 39.0 | 51.5 | 38.5 | 41.5 | 18.0 | 23.0 | 17.3 | 13.0 | 28.1 | 31.0 | ||||
Llama-3 | 8B | 49.5 | 56.5 | 21.5 | 31.0 | 24.0 | 29.0 | 10.0 | 12.3 | 24.5 | 30.1 | |||||
Qwen2 | 7B | 13.0 | 56.0 | 9.5 | 33.0 | 4.0 | 31.0 | 2.3 | 10.0 | 7.0 | 29.9 | |||||
Mathstral | 7B | Math | 43.5 | 55.0 | 32.5 | 35.0 | 10.0 | 23.0 | 11.3 | 11.7 | 24.5 | 29.8 | ||||
GLM-4 | 9B | 69.5 | 44.0 | 53.5 | 34.0 | 33.0 | 20.0 | 17.7 | 8.7 | 41.5 | 25.3 | |||||
Aya-23 | 35B | 1.5 | 44.0 | 1.0 | 25.5 | 0.0 | 20.0 | 0.0 | 11.7 | 0.6 | 24.3 | |||||
DeepSeek-V2-Lite | 16B | MoE | 7.0 | 45.5 | 3.5 | 18.0 | 1.0 | 17.0 | 1.0 | 10.3 | 3.1 | 21.9 | ||||
Mixtral-8x7B-v0.1 | 46B | MoE | 0.5 | 39.0 | 2.0 | 17.0 | 0.0 | 25.0 | 0.0 | 12.7 | 0.6 | 21.9 | ||||
DeepSeek-Math | 7B | Math | 2.0 | 46.0 | 1.0 | 27.0 | 1.0 | 4.0 | 0.3 | 8.0 | 1.0 | 21.8 | ||||
Llama-2 | 70B | 32.5 | 43.5 | 16.5 | 25.0 | 1.0 | 8.0 | 2.0 | 7.0 | 13.1 | 20.8 | |||||
WizardLM-2 | 7B | 47.0 | 42.0 | 30.5 | 28.5 | 5.0 | 6.0 | 7.3 | 5.7 | 22.7 | 20.5 | |||||
Mistral-v0.3 | 7B | 49.5 | 40.0 | 40.5 | 28.0 | 25.0 | 9.0 | 11.3 | 5.7 | 29.9 | 20.3 | |||||
WizardMath | 7B | Math | 22.5 | 32.0 | 12.0 | 22.5 | 6.0 | 7.0 | 3.7 | 3.3 | 10.8 | 15.7 | ||||
InternLM2-Math-Plus | 7B | Math | 28.5 | 27.5 | 15.0 | 14.0 | 7.0 | 9.0 | 4.7 | 4.0 | 13.5 | 13.0 | ||||
StarCoder2 | 15B | Code | 47.5 | 21.0 | 34.0 | 15.5 | 11.0 | 6.0 | 8.3 | 4.3 | 24.9 | 11.5 | ||||
InternLM2 | 7B | 18.0 | 20.0 | 4.5 | 11.0 | 9.0 | 10.0 | 2.7 | 2.3 | 7.8 | 9.9 | |||||
Gemma-1 | 7B | 1.0 | 20.0 | 0.0 | 7.5 | 0.0 | 7.0 | 0.0 | 3.3 | 0.2 | 9.0 | |||||
Llama-2 | 7B | 4.0 | 17.0 | 4.0 | 11.5 | 0.0 | 2.0 | 1.3 | 2.7 | 2.5 | 8.4 | |||||
DeepSeek-Coder-V1 | 33B | Code | 19.0 | 18.5 | 8.5 | 8.5 | 2.0 | 2.0 | 3.7 | 1.7 | 8.5 | 7.6 | ||||
WizardCoder | 33B | Code | 32.5 | 16.0 | 17.5 | 8.0 | 5.0 | 2.0 | 5.0 | 1.0 | 15.0 | 6.6 | ||||
Aya-23 | 8B | 1.0 | 13.0 | 0.0 | 9.0 | 0.0 | 2.0 | 0.3 | 2.3 | 0.4 | 6.6 | |||||
Gemma-1 | 2B | 4.0 | 8.0 | 1.5 | 7.5 | 0.0 | 2.0 | 0.0 | 0.0 | 1.4 | 4.1 |
4.3 Implementation Details
LLM Experiment
The experiments for open-sourced LLMs were conducted using the vLLM framework Kwon et al. (2023). For all the experiments, we set temperature as 1.0 and maximum output length as 512. Given the extensive context length of input document, the main evaluation of DocMath-Eval is conducted under a zero-shot setting, aiming to assess LLMs’ capabilities to generate accurate answers without few-shot demonstrations or additional training.
Input Tabular Data Serialization
Building on previous work that evaluated LLMs on table-relevant tasks Chen (2023); Zhao et al. (2023b, c), we present our method for processing tabular data in documents. Specifically, we separate headers or cells in different columns using a vertical bar (|), and rows using a newline. This approach allows for the direct feeding of flattened table input into LLMs. In our preliminary study, we discovered that most LLMs can effectively understand such table representations. Nevertheless, we believe future research could explore more effective methods for encoding tabular data.
RAG-based Setting for DM
For the DM subset, the input document length is extremely long and exceeds the context length limit of evaluated LLMs. Therefore, in our main experiments with DM, we evaluate models using the retrieval-augmented generation (RAG) setting. In this setting, external retrievers are employed to extract the top- most relevant textual and tabular evidence from the source document. We maintain the original relative order of the evidence and input it into the LLMs to answer the given question. We experiment with commonly-used (1) sparse retriever, i.e., BM25 Robertson et al. (1995), and (2) three dense retrievers, including OpenAI Embedding 3 small & large versions Neelakantan et al. (2022) and Contriever Izacard et al. (2022).
Final Answer Extraction
For LLM with CoT prompting, we adopt the answer extraction pipeline from Chen et al. (2023b) to identify the final answer from the model’s output. For LLM with PoT prompting, we first develop a heuristic method to extract the generated python solution from the model’s output. If this python solution is executable, we execute it to obtain the final answer.
5 Results and Analysis
5.1 Main Results
Table 2 and Table 9 in the Appendix present the LLM performance using CoT and PoT prompting methods on the DocMath-Eval testmini and test sets, respectively. While the current best-performing LLM, GPT-4o, achieves performance comparable to human experts in simple problem settings (i.e., DM and DM), we find significant performance gaps in more challenging settings. Specifically, GPT-4o achieves an accuracy of 41.0% on DM with PoT, which is far behind the human expert performance of 76.0%. This underscores the need for ongoing LLM development, particularly in complex problem-solving over long and specialized documents.
Most open-source LLMs still lag behind the proprietary LLMs. However, the two DeepSeek-V2 models achieve performance levels close to the best-performing proprietary models. This indicates the potential for open-source LLMs to close the performance gap with proprietary models in the near future, given continued innovation and community collaboration. Additionally, the proprietary LLMs and code-specific models typically achieve comparable or better performance when using PoT prompting compared to CoT prompting. For math-specific LLMs, InternLM2-Math-Plus surpasses its backbone models in CoT performance, improving from 9.9% to 13.0% on average accuracy. This demonstrates the effectiveness of instruction-tuning in enhancing math reasoning.
5.2 Analysis on DM Set
We next conduct a detailed analysis of the RAG setting, long-context LLMs, and model failure cases.
top- | Retriever | R@ | Llama-3 | GPT-4o |
3 | Contriever | 22.3 | 13.7 | 16.0 |
BM25 | 29.5 | 13.7 | 15.7 | |
Embedding-3-Small | 44.7 | 19.0 | 24.0 | |
Embedding-3-Large | 48.2 | 22.0 | 27.0 | |
[…]
BM25 | 38.0 | 15.3 | 20.7 | |
Embedding-3-Small | 57.1 | 21.0 | 29.0 | |
Embedding-3-Large | 62.0 | 24.0 | 32.7 | |
[…]
BM25 | 47.9 | 20.3 | 23.0 | |
Embedding-3-Small | 71.2 | 25.3 | 31.7 | |
Embedding-3-Large | 75.8 | 26.3 | 36.7 | |
– | Oracle | 35.3 | 42.0 |
RAG Analysis
We analyze the impact of retriever performance on the final accuracy of RAG-based LLM systems by selecting the Llama-3-70B and GPT-4o models for our study. As demonstrated in subsection 5.2, the OpenAI Embedding-3 significantly outperforms Contriever and BM25. Additionally, improved retriever performance consistently boosts the final accuracy of the models in our task. These results highlight the need for future work to develop more advanced information retrieval techniques for enhancing complex problem-solving over long and specialized documents.
Model | RAG | Long Context |
GPT-4o | 36.7 | 40.3 |
DeepSeek-V2 | 39.7 | 38.7 |
DeepSeek-Coder-V2 | 37.7 | 36.0 |
Gemini-1.5-Pro | 30.0 | 37.3 |
Claude-3-Sonnet | 35.3 | 34.7 |
Gemini-1.5-Flash | 31.7 | 34.3 |
Claude-3-Haiku | 31.7 | 31.0 |
Llama-3.1-70B | 29.3 | 26.3 |
Llama-3.1-8B | 14.3 | 9.0 |
Mistral-Nemo | 16.7 | 4.7 |
Phi-3-Medium | 12.7 | 13.0 |
GLM-4-9B | 8.7 | 9.7 |
Long-Context LLM Analysis
In addition to using RAG for analyzing long specialized documents, recent advancements have extended the input length of LLMs to handle lengthy documents Su et al. (2023). We compare models with a context length limit of over 100K under both the RAG (as used in the main results) and Long-Context settings, where the entire document is input. As illustrated in Table 4, the evaluated models generally achieve close performance under RAG and long-context settings. This indicates that models with extended context lengths can effectively process lengthy inputs without a significant drop in performance compared to the RAG setting.
Error Analysis
To better understand the strengths and weaknesses of LLMs, we conduct an extensive error analysis. This analysis focuses on 100 randomly selected examples from the DM testmini set where GPT-3.5-turbo failed. We identify four common types of errors in current LLMs: inaccurate evidence retrieval, calculation errors, table misunderstandings, and exceeding context length. A detailed explanation for each type is provided in Appendix A in Appendix.
6 Conclusion
This paper introduces DocMath-Eval, a comprehensive benchmark designed to evaluate the capabilities of LLMs in numerical reasoning over long and specialized documents. Our experiments show that even the best-performing current models still fall short of human expert performance on problems requiring complex reasoning over extended contexts. This highlights the need for future research to improve LLMs’ proficiency in complex numerical reasoning tasks within expert domains.
Limitations
In this work, we propose DocMath-Eval and conduct comprehensive analysis of different LLMs’ capabilities in solving knowledge-intensive math reasoning problems in finance domains. However, there are still some limitations: First, our method for extracting final answer from model output is still not perfect. In some cases, this methods fails to locate the answer, leading to the reported accuracy being an approximate lower bound. Moreover, due to computational resource constraints, we do not tune LLMs on a large-scale finance-domain data ourselves. However, we believe that training on finance data can help improve LLMs’ capabilities in solving problems in DocMath-Eval Wu et al. (2023b); Luukkonen et al. (2023); Xie et al. (2023).
Acknowledgement
We are grateful for the compute support provided by Microsoft Research’s Accelerate Foundation Models Research (AFMR) program. We would also like to thank the anonymous reviewers and area chairs for constructive discussions and feedback.
Appendix A Appendix
Annotation Quality | %S 4 |
Question Fluency | 97.4 |
Question Correctness | 96.0 |
[…]
Evidence Completeness | 91.3 |
[…]
Python Solution Correctness | 97.6 |
Variable Value Correctness | 98.5 |
Python Solution Conciseness | 89.1 |
Variable Name Meaningfulness | 95.4 |
Comment Comprehensiveness | 87.4 |
Error Type | Explanation |
Inaccurate Evidence Retrieval
(39 / 100) |
The challenge lies in finding accurate evidence, especially in situations where the values needed for intermediate reasoning steps are not explicitly stated. This makes it difficult for the retriever to identify the correct evidence. |
[…]
Calculation Error
(28 / 100) |
The reasoning process is accurate, but there are errors in the intermediate or final computations. |
[…]
Table Misunderstanding
(16 / 100) |
The model faces challenges in comprehending and parsing cell values, particularly in complex tables. |
[…]
Exceeding Context Length
(8 / 100) |
The input document exceeds the context length limit. |
[…]
Annotator ID | Finance Industry Experience | Annotation Sets |
1 | 1 working and 1 internship at US | New subset, Annotation validation |
2 | >= 2 internship at US | New subset, Annotation validation |
3 | 1 working at Singapore and 2 internship at US | New subset |
4 | 2 working and >= 1 internship at US | New subset |
5 | 1 internship at US, 2 internship at China | Re-annotation on three subsets, Annotation validation |
6 | Graduate student majored in computer science | Re-annotation on three subsets, Annotation validation |
7 | Graduate student majored in statistics | Re-annotation on three subsets |
Model | Organization | Size | Notes | Source |
GPT-4-Turbo | OpenAI | – | gpt-4o-2024-05-13 | |
GPT-4o | OpenAI | – | gpt-4-turbo-2024-04-09 | |
GPT-3.5-Turbo | OpenAI | – | gpt-3.5-turbo-0125 | |
[…]
Claude-3-Opus | Anthropic | – | claude-3-opus-20240229 | |
Claude-3-Sonnet | Anthropic | – | claude-3-sonnet-20240229 | |
Claude-3-Haiku | Anthropic | – | claude-3-haiku-20240307 | |
[…]
Gemini-1.5-Flash | – | gemini-1.5-flash | ||
Qwen2 | Alibaba | 7 & 72B | Qwen/Qwen2-*B-Instruct | |
[…]
Llama-3 | Meta | 8 & 70B | meta-llama/Meta-Llama-3-*B-Instruct | |
Llama-3.1 | Meta | 8 & 70B & 405B | meta-llama/Meta-Llama-3.1-*B-Instruct | |
[…]
Gemma-2 | 9B | google/gemma-2-9b-it | ||
[…]
Mistral-Nemo | Mistral AI | 12B | mistralai/Mistral-Nemo-Instruct-2407 | |
Mistral-Large | Mistral AI | 123B | mistralai/Mistral-Large-Instruct-2407 | |
Mathstral | Mistral AI | 7B | Math-Specific | mistralai/Mathstral-7B-v0.1 |
Mixtral | Mistral AI | 46 & 141B | MoE | mistralai/Mixtral--Instruct-v0.1 |
Codestral | Mistral AI | 22B | Code-Specific | mistralai/Codestral-22B-v0.1 |
[…]
DeepSeek-Coder-V1 | DeepSeek | 33B | Code-Specific | deepseek-ai/deepseek-coder-33b-instruct |
DeepSeek-V2 | DeepSeek | 16 & 236B | MoE | deepseek-ai/DeepSeek-V2-Lite-Chat. We use the official API provided by DeepSeek for deepseek-ai/DeepSeek-V2-Chat |
DeepSeek-Coder-V2 | DeepSeek | 16 & 236B | MoE | deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct. We use the official API provided by DeepSeek for deepseek-ai/DeepSeek-Coder-V2-Instruct |
[…]
Phi-3-Mini | Microsoft | 3B | microsoft/Phi-3-mini-4k-instruct | |
[…]
InternLM2-Math-Plus | InternLM | 7B | Math-Specific | internlm/internlm2-math-plus-7b |
[…]
WizardMath | WizardLM Team | 7B | Math-Specific | WizardLMTeam/WizardMath-7B-V1.1 |
WizardCoder | WizardLM Team | 33B | Code-Specific | WizardLMTeam/WizardCoder-33B-V1.1 |
WizardLM-2 (MoE) | WizardLM Team | 141B | MoE | alpindale/WizardLM-2-8x22B |
[…]
Model | Size | Notes | DM | DM | DM | DM | Avg. Acc | |||||||||
PoT | CoT | PoT | CoT | PoT | CoT | PoT | CoT | PoT | CoT | |||||||
Human Expert | 91.0 | 87.0 | 84.0 | 76.0 | ||||||||||||
Proprietary LLMs | ||||||||||||||||
GPT-4-Turbo | 88.9 | 86.2 | 78.6 | 77.8 | 31.5 | 61.2 | 20.2 | 35.6 | 53.4 | 62.0 | ||||||
GPT-4o | 87.0 | 86.4 | 69.6 | 75.9 | 62.2 | 66.2 | 41.5 | 31.7 | 62.5 | 60.7 | ||||||
Claude-3.5-Sonnet | 87.2 | 81.8 | 71.0 | 70.0 | 65.0 | 61.5 | 40.8 | 31.5 | 63.0 | 57.4 | ||||||
GPT-4o-Mini | 88.8 | 76.0 | 77.2 | 72.4 | 56.8 | 50.2 | 36.9 | 29.2 | 62.4 | 54.3 | ||||||
Gemini-1.5-Pro | 87.9 | 82.5 | 80.4 | 58.4 | 63.0 | 58.0 | 41.8 | 31.2 | 65.6 | 54.2 | ||||||
Claude-3-Haiku | 77.8 | 82.2 | 72.4 | 61.5 | 55.0 | 53.2 | 34.6 | 30.8 | 57.4 | 54.2 | ||||||
Gemini-1.5-Flash | 87.9 | 80.5 | 79.9 | 67.0 | 60.5 | 53.8 | 38.9 | 28.1 | 64.1 | 54.1 | ||||||
Claude-3-Sonnet | 86.9 | 84.9 | 80.0 | 69.0 | 57.8 | 54.5 | -1.0 | -1.0 | 48.9 | 45.3 | ||||||
GPT-3.5-Turbo | 75.6 | 64.0 | 47.6 | 34.2 | 46.2 | 38.8 | 23.2 | 13.2 | 45.3 | 34.3 | ||||||
Open-source LLMs | ||||||||||||||||
Mistral-Large | 123B | 87.9 | 85.0 | 74.4 | 79.8 | 58.8 | 58.8 | 37.2 | 30.6 | 61.8 | 60.0 | |||||
DeepSeek-V2 | 236B | MoE | 88.9 | 86.1 | 76.9 | 66.6 | 58.2 | 57.2 | 40.2 | 32.0 | 63.8 | 57.3 | ||||
DeepSeek-Coder-V2 | 236B | Code | 88.2 | 80.9 | 74.9 | 67.2 | 56.8 | 54.8 | 37.5 | 31.5 | 61.9 | 55.7 | ||||
Llama-3.1 | 70B | 81.0 | 80.1 | 65.8 | 70.8 | 49.2 | 53.8 | 32.1 | 26.6 | 54.9 | 54.4 | |||||
Qwen2 | 72B | 27.8 | 77.5 | 25.0 | 70.1 | 16.8 | 49.0 | 5.7 | 25.1 | 17.4 | 52.4 | |||||
Llama-3 | 70B | 86.2 | 80.9 | 64.8 | 62.0 | 51.5 | 45.0 | 35.0 | 26.3 | 57.3 | 51.2 | |||||
Mixtral-8x22B | 141B | MoE | 27.4 | 74.9 | 23.1 | 59.4 | 21.5 | 46.8 | 14.8 | 22.3 | 20.9 | 47.8 | ||||
DeepSeek-Coder-V2-Lite | 16B | Code | 68.8 | 71.4 | 52.2 | 44.9 | 28.8 | 35.0 | 18.0 | 19.8 | 40.6 | 40.9 | ||||
Gemma-2 | 9B | 82.0 | 70.1 | 62.4 | 49.2 | 45.5 | 33.2 | 24.4 | 17.7 | 50.9 | 40.6 | |||||
Yi-1.5 | 34B | 1.0 | 71.2 | 0.9 | 58.2 | 0.2 | 17.8 | 0.2 | 13.7 | 0.6 | 39.7 | |||||
C4AI Command R+ | 104B | 37.6 | 67.5 | 36.5 | 50.6 | 19.2 | 35.0 | 7.6 | 15.1 | 23.8 | 39.6 | |||||
WizardLM-2 | 141B | MoE | 62.1 | 59.4 | 49.5 | 47.6 | 31.2 | 36.0 | 19.0 | 16.1 | 38.9 | 37.3 | ||||
DBRX | 132B | MoE | 46.8 | 64.0 | 33.1 | 41.5 | 26.5 | 29.8 | 10.7 | 18.4 | 27.3 | 37.0 | ||||
Mistral-Nemo | 12B | 51.1 | 66.0 | 32.1 | 40.5 | 29.0 | 37.0 | 15.7 | 15.0 | 30.3 | 36.9 | |||||
Yi-1.5 | 9B | 23.2 | 69.5 | 17.4 | 45.0 | 1.2 | 15.0 | 2.2 | 11.9 | 11.1 | 35.0 | |||||
Llama-3.1 | 8B | 66.5 | 63.0 | 41.9 | 34.1 | 35.5 | 35.5 | 14.7 | 13.8 | 37.0 | 33.9 | |||||
Codestral | 22B | Code | 43.1 | 58.0 | 36.8 | 37.2 | 25.0 | 29.5 | 14.2 | 14.0 | 28.4 | 32.8 | ||||
Llama-3 | 8B | 51.6 | 57.8 | 22.8 | 31.2 | 22.5 | 25.0 | 9.3 | 11.5 | 24.9 | 29.7 | |||||
Mathstral | 7B | Math | 45.8 | 54.4 | 31.1 | 34.0 | 11.0 | 24.0 | 9.1 | 11.7 | 24.0 | 29.5 | ||||
Qwen2 | 7B | 15.4 | 52.6 | 6.6 | 33.2 | 4.2 | 29.0 | 2.6 | 11.4 | 7.0 | 29.4 | |||||
GLM-4 | 9B | 68.0 | 48.8 | 46.5 | 32.1 | 32.5 | 22.8 | 18.2 | 11.1 | 39.5 | 27.2 | |||||
Aya-23 | 35B | 0.9 | 46.9 | 0.5 | 26.5 | 0.0 | 19.5 | 0.5 | 10.1 | 0.5 | 24.6 | |||||
Mixtral-8x7B-v0.1 | 46B | MoE | 1.1 | 42.5 | 0.4 | 19.6 | 0.2 | 24.2 | 0.2 | 13.0 | 0.5 | 23.4 | ||||
DeepSeek-Math | 7B | Math | 1.4 | 47.1 | 0.4 | 28.0 | 1.2 | 12.5 | 0.5 | 7.9 | 0.8 | 23.3 | ||||
Mistral-v0.3 | 7B | 48.6 | 40.8 | 28.5 | 24.8 | 19.5 | 18.0 | 12.8 | 7.6 | 26.5 | 21.5 | |||||
DeepSeek-V2-Lite | 16B | MoE | 7.6 | 49.1 | 4.6 | 16.9 | 2.5 | 15.2 | 1.0 | 8.3 | 3.7 | 21.5 | ||||
Llama-2 | 70B | 27.1 | 45.1 | 16.8 | 26.0 | 2.0 | 8.0 | 1.1 | 6.9 | 11.6 | 21.4 | |||||
WizardLM-2 | 7B | 47.0 | 42.6 | 31.2 | 31.2 | 9.2 | 8.5 | 7.2 | 4.8 | 23.4 | 21.3 | |||||
WizardMath | 7B | Math | 22.1 | 34.2 | 14.4 | 24.8 | 5.8 | 6.2 | 3.4 | 4.6 | 11.1 | 17.2 | ||||
InternLM2-Math-Plus | 7B | Math | 30.0 | 30.2 | 15.0 | 16.2 | 10.8 | 11.2 | 4.2 | 3.3 | 14.2 | 14.3 | ||||
StarCoder2 | 15B | Code | 51.0 | 29.8 | 32.0 | 16.8 | 9.5 | 5.2 | 8.4 | 3.0 | 25.1 | 13.4 | ||||
InternLM2 | 7B | 17.4 | 24.5 | 9.4 | 11.8 | 9.0 | 6.8 | 2.9 | 4.3 | 8.9 | 11.5 | |||||
Gemma-1 | 7B | 0.5 | 23.2 | 0.2 | 6.0 | 0.0 | 7.8 | 0.2 | 3.2 | 0.3 | 9.5 | |||||
Llama-2 | 7B | 5.6 | 20.4 | 2.4 | 11.8 | 1.0 | 3.2 | 0.4 | 2.2 | 2.3 | 9.3 | |||||
WizardCoder | 33B | Code | 38.4 | 19.9 | 17.5 | 8.4 | 7.2 | 3.2 | 6.1 | 1.8 | 17.2 | 8.2 | ||||
Aya-23 | 8B | 0.5 | 14.2 | 0.1 | 8.5 | 0.0 | 2.8 | 0.0 | 2.2 | 0.2 | 6.8 | |||||
DeepSeek-Coder-V1 | 33B | Code | 19.4 | 16.2 | 8.5 | 6.6 | 3.2 | 2.2 | 3.6 | 1.2 | 8.7 | 6.4 | ||||
Gemma-1 | 2B | 5.9 | 9.1 | 2.5 | 6.0 | 0.8 | 2.8 | 0.2 | 1.0 | 2.3 | 4.5 |